Compute the test statistic used to test the null hypothesis that p1 = p2.
In a vote on the Clean Water bill, 53% of the 205 Democrats voted for the bill while 51% of the 230 Republicans voted for it.

Answer by stanbon(48568)

Compute the test statistic used to test the null hypothesis that p1 = p2.
In a vote on the Clean Water bill, 53% of the 205 Democrats voted for the bill while 51% of the 230 Republicans voted for it.
---------------------------------------------------------
Ho: p1-p2 = 0
----
ts = (0.53-0.51)/sqrt[(0.53*0.47)/205 + (0.51*0.49/230)]
= 0.02/0.0480
= 0.4169
==============
Cheers,
Stan H.